1. The stock of Blue Water Tours, Inc. is expected to return 21.50 percent in a...

1.

The stock of Blue Water Tours, Inc. is expected to return 21.50 percent in a boom economy, 16.50 percent in a normal economy, and lose 15.50 percent in a recessionary economy. What is the expected rate of return on this stock if there is a 7.00 percent chance the economy booms, and an 83.00 percent chance the economy will be normal?

14.13 percent

13.65 percent

13.40 percent

12.48 percent

2.

A stock is expected to earn 15 percent in a boom economy and 7 percent in a normal economy. There is a 35 percent chance the economy will boom and a 65.0 percent chance the economy will be normal. What is the standard deviation of these returns?

3.82 Percent

4.85 Percent

4.97 Percent

5.63 Percent

3.

A portfolio consists of 24 percent Stock A, 54 percent Stock B, and 22 percent Stock C. What is the portfolio expected return given the following:

State of Economy

Probability of State of Economy

Stock A Returns

Stock B Returns

Stock C Returns

Normal

.75

16%

   9%

   26%

Recession

.25

–2    

19

–24

10.55 percent

11.94 percent

8.24 percent

12.47 percent
Answer #1

(1) Return in case of :

Boom = 21.5 % , Normal = 16.5 % and Recession = - 15.5 %

Probability of:

Boom = 7 %, Normal = 83 % and Recession = (100 - 7 - 83) = 10 %

Expected Return = 21.5 x 0.07 + 16.5 x 0.83 - 15.5 x 0.1 = 13.65 %

Hence, the correct option is (b).

NOTE: Please raise separate queries for solutions to the remaining unrelated questions as one query is restricted to the solution of only one distinct question.